# A Variational Framework for the Algorithmic Complexity of PDE Solutions

### v1

@@ 7.2 Complexity of Solutions to the Eikonal Equation
This subsection presents the Eikonal equation [15] as a representative example of a PDE whose solution is $H^{1}$-computable but exhibits a loss of regularity, resulting in a complexity blowup. We begin by formulating the corresponding variational problem.
@@
The Eikonal loss $\mathcal{L}:H^{1}(\Omega)\rightarrow\mathbb{R}_{+}$ from (68) is a convex, continuous functional, satisfying the apriori coercivity estimate
$\|u\|_{H^{1}(\Omega)}\lesssim\mathcal{L}[u]+1,$ (69)
and the QGC. Therefore, there exists a solution $u^{*}\in H^{1}(\Omega)$ minimizing $\mathcal{L}$.
@@
We start by proving the convexity of the loss. Let $u,v\in H^{1}(\Omega)$ and $t\in(0,1)$, then by the monotonicity and convexity of the norms and the linearity of the gradient, we get
$\mathcal{L}[tu+(1-t)v]$ $\leq\bigr\|t\|\nabla u\|_{1}+(1-t)\|\nabla v\|_{1}-1\bigr\|_{L^{2}(\Omega)}+t\|u|_{S}\|_{H^{1/2}(S)}+(1-t)\|v|_{S}\bigr\|_{H^{1/2}(S)}$
$\leq t\mathcal{L}[u]+(1-t)\mathcal{L}[v].$
Next, we show the Hölder continuity of the loss. Using the reverse triangle inequality, the following holds
@@
It remains to prove the quadratic growth condition of the functional. Let $u\in H^{1}(\Omega)$, and $u^{*}\in H^{1}(\Omega)$ the minimizer of the loss $\mathcal{L}$. Applying the reverse triangle inequality, the estimate in (70), and assuming without loss of generality that $\mathcal{L}[u^{*}]=0$, the following holds
$\mathcal{L}[u]-\mathcal{L}[u^{*}]$ $=\bigr\|\|\nabla u\|_{1}-1\bigr\|_{L^{2}(\Omega)}+\|u|_{S}\|_{H^{1/2}(S)}$ (75)
$\geq\bigr\|\|\nabla u-\nabla u^{*}\|_{1}-\|\nabla u^{*}\|_{1}-1\bigr\|_{L^{2}(\Omega)}+\|u|_{S}\|_{H^{1/2}(S)}$ (76)
$=\bigr\|\|\nabla u-\nabla u^{*}\|_{1}\bigr\|_{L^{2}(\Omega)}+\|u|_{S}\|_{H^{1/2}(S)}$ (77)
$=\bigr\|\|\nabla u-\nabla u^{*}\|_{1}\bigr\|_{L^{2}(\Omega)}+\|u|_{S}-u^{*}|_{S}\|_{H^{1/2}(S)}$ (78)
$\geq\bigr\|\nabla u-\nabla u^{*}\bigr\|_{L^{2}(\Omega)}+\|u|_{S}-u^{*}|_{S}\|_{H^{1/2}(S)}$ (79)
$\geq\frac{1}{2}\|u-u^{*}\|_{H^{1}(\Omega)}.$ (80)
For the third inequality we used that since $\mathcal{L}[u^{*}]=0$, then $\|\nabla u^{*}\|_{1}-1=0$ almost everywhere in $\Omega$. ∎
@@
This is a well-known result in the signed distance function literature [40], and we include a proof in Appendix C. ∎
@@
We start by showing the existence of the solution $u^{*}\in H^{1}(\Omega)$. By Theorem 7.7 we know that the Eikonal equation loss $\mathcal{L}:H^{1}(\Omega)\rightarrow\mathbb{R}_{+}$ from (68) is continuous, convex and coercive, hence by Tonelli’s direct method, there exists a solution $u^{*}\in H^{1}(\Omega)$. Moreover, by Theorem 7.7 we know that $\mathcal{L}:H^{1}(\Omega)\rightarrow\mathbb{R}_{+}$ satisfies the quadratic growth condition.
Next, we note that the operator $\mathcal{N}:H^{1}(\Omega)\rightarrow L^{2}(\Omega)$, $\mathcal{N}[\hat{u}_{\theta}]:=\|\nabla\hat{u}_{\theta}\|_{1}$ is $C^{0}$-computable for all $\hat{u}_{\theta}\in\Pi_{n}(\Omega)$. In general, the gradient $\nabla\mathcal{L}^{n}$ is not globally Lipschitz continuous, since it fails to be Lipschitz in regions where $\|\nabla\hat{u}_{\theta^{j}}\|_{1}$ approaches zero. Consequently, Theorem 6.13 cannot be applied directly, as the convergence rates for the explicit Euler gradient flow no longer hold. Instead, we can consider a more general framework, such as the Euler sub-gradient flow [41], which requires only local Lipschitz continuity and yields convergence rates of order $\mathcal{O}(\sqrt{j})$.
To prove the complexity blowup of the solution, we assume that $S\subseteq\Omega$ is given by a polynomial time, $H^{1}$-computable parametrization $\Psi:U\rightarrow\Omega$, for some open set $U\subseteq\mathbb{R}^{d}$. Following [15], we use the well-known result that the signed distance function $\phi\in H^{1}(\Omega)$, defined in (81), is a minimizer of the loss functional $\mathcal{L}$; in particular, it is the viscosity solution of the Eikonal equation [42]. By Lemma 7.9, there exists a subset $M_{s}\subseteq\{\phi\geq 0\}$ (the medial axis of the manifold $S$) where $\phi$ is not differentiable, and therefore the function $\phi$ is not $\rho$-analytic. It follows from Corollary 6.16 that the solution $\phi$ is $H^{1}$-computable in at least super-polynomial time, establishing the claim. ∎
@@ 8 Conclusion
In this paper, we introduced a variational framework for analyzing the computational complexity of partial differential equations (PDEs) without requiring closed-form solutions, thereby generalizing results from previous works [5, 17]. Specifically, the variational framework allows to estimate the complexity of a discrete gradient flow, approximation the solution operator of the variational problem, and uses the convergence rates of these discretizations to estimate the complexity of the infinite-dimensional solution of the PDE. This approach enabled us to relate structural properties of PDE operators and their associated variational losses to the computational complexity of the corresponding solutions. Moreover, we introduce a method for proving complexity blowup of the solution based on the regularity-preserving properties of the solution operators. Specifically, if the solution loses regularity relative to analytical boundary and/or initial data, then the solution exhibits a complexity blowup. This result gives a direct link between regularity and approximation theory of PDEs and complexity theory.
@@ A Banach space $X$ is said to have a computability structure, if the following axioms are fulfilled:
One way to prove the polynomial computability of the Legendre grid and weights is to use the Golub–Welsch (GW) algorithm [43]. For simplicity, consider the case $d=1$. Let $J_{n}\in\mathbb{R}^{(n+1)\times(n+1)}$ be the Jacobi matrix, defined as the tri-diagonal matrix with entries $(J_{n})_{i,j}=0$ for $j\not\in\{i+1,i\}$ for $i\in\{1,\dots,n\}$ and $(J_{n})_{i-1,i}=\beta_{i+1}$$(J_{n})_{i+1,i}=\beta_{i}$ for $i\in\{2,\dots,n\}$, where $\beta_{i}:=\frac{i}{\sqrt{(2i-1)(2i+1)}}$. The Legendre nodes $P_{n}(\Omega)$ are then obtained as the eigenvalues of $J_{n}$. Since the coefficients $\beta_{i}$ are computable in polynomial time, and the tri-diagonal structure allows eigenvalues of $J_{n}$ to be computed in $\mathcal{O}(n^{2})$ operations [44], this yields the desired complexity bound.
An alternative proof of polynomial computability uses more modern approaches, such as Newton–Raphson iterations with carefully chosen initializations [45]. These methods achieve improved complexity, requiring only $\mathcal{O}(n)$ [44] or even $\mathcal{O}(1)$ [46] operations, and currently represent the state of the art for computing Legendre quadratures. ∎
@@
First, observe that the Legendre polynomials can be expressed via the Rodriguez formula [47] as derivatives of the polynomial $(x^{2}-1)^{n}$, i.e., the 1-D Legendre polynomial of degree $n\in\mathbb{N}$ is given by
@@
Since $(x^{2}-1)^{n}$ is analytic and $C^{0}$-computable, its derivatives are also $C^{0}$-computable [31] and therefore $H^{k}$-computable. In higher dimensions, the Legendre polynomial $P_{\alpha}\in\Pi_{n}(\Omega)$ for $\alpha\in A_{n,d}$ is given as a finite product of 1-D basis elements. Consequently, the Legendre polynomials themselves are $C^{0}$-computable and therefore $H^{k}$-computable.
Moreover, the roots of these polynomials are computable [31], ensuring that the Legendre grid $P_{n}(\Omega)\subseteq\Omega$ is computable. By the $H^{k}$-computability of $f$, the finite sequence $f(P_{n}(\Omega)):=(f(p_{\alpha}))_{\alpha\in A_{n,d}}\subseteq\mathbb{R}$ is also computable.
@@ Appendix C Appendix
Next, we show the coercivity estimate from (58). Let $u\in H^{1}(\Omega)$ and $\mathcal{E}:H^{1/2}(\partial\Omega)\rightarrow H^{1}(\Omega)$ be the extension operator [48]. We define $u_{b}:=\mathcal{E}[u|_{\partial\Omega}]$ and $u_{0}:=u-u_{b}$. Then, by Cauchy-Schwarz, the triangle inequality and the extension theorem [49], the following holds
@@
Therefore, applying the estimates in (105) and 101, we have that
@@
To finish the proof we apply Tonelli’s direct method [50] to a convex, coercive and lower-semi-continuous functional, yielding the existence of the solution. The uniqueness holds since $\mathcal{L}$ is $\lambda$-convex and therefore strictly convex. ∎

### v2

@@ 7.2 Complexity of Solutions to the Eikonal Equation
This subsection presents the Eikonal equation [Alimov2019-np] as a representative example of a PDE whose solution is $H^{1}$-computable but exhibits a loss of regularity, resulting in a complexity blowup. We begin by formulating the corresponding variational problem.
@@
The Eikonal loss $\mathcal{L}:H^{1}(\Omega)\rightarrow\mathbb{R}_{+}$ from (68) is a continuous functional, satisfying the apriori coercivity estimate
$\|u\|_{H^{1}(\Omega)}\lesssim\mathcal{L}[u]+1.$ (69)
Therefore, there exists a solution $u^{*}\in H^{1}(\Omega)$ minimizing $\mathcal{L}$.
@@
We start by proving the continuity of the loss. Using the reverse triangle inequality, the following holds
@@
This is a well-known result in the signed distance function literature [Alimov2019-gh], and we include a proof in Appendix C. ∎
@@
We start by showing the existence of the solution $u^{*}\in H^{1}(\Omega)$. By Theorem 7.7 we know that the Eikonal equation loss $\mathcal{L}:H^{1}(\Omega)\rightarrow\mathbb{R}_{+}$ from (68) is continuous, and coercive, hence by Tonelli’s direct method, there exists a solution $u^{*}\in H^{1}(\Omega)$.
Next, we note that the operator $\mathcal{N}:H^{1}(\Omega)\rightarrow L^{2}(\Omega)$, $\mathcal{N}[\hat{u}_{\theta}]:=\|\nabla\hat{u}_{\theta}\|_{1}$ is $C^{0}$-computable for all $\hat{u}_{\theta}\in\Pi_{n}(\Omega)$. In general, the gradient $\nabla\mathcal{L}^{n}$ is not globally Lipschitz continuous, since it fails to be Lipschitz in regions where $\|\nabla\hat{u}_{\theta^{j}}\|_{1}$ approaches zero. Consequently, Theorem 6.13 cannot be applied directly, as the convergence rates for the explicit Euler gradient flow no longer hold. Instead, we can consider a more general framework, such as the Euler sub-gradient flow [Li2023-ud], which requires only local Lipschitz continuity and yields convergence rates of order $\mathcal{O}(\sqrt{j})$.
To prove the complexity blowup of the solution, we assume that $S\subseteq\Omega$ is given by a polynomial time, $H^{1}$-computable parametrization $\Psi:U\rightarrow\Omega$, for some open set $U\subseteq\mathbb{R}^{d}$. Following [Alimov2019-np], we use the well-known result that the signed distance function $\phi\in H^{1}(\Omega)$, defined in (75), is a minimizer of the loss functional $\mathcal{L}$; in particular, it is the viscosity solution of the Eikonal equation [Katzourakis2014-fe]. By Lemma 7.9, there exists a subset $M_{s}\subseteq\{\phi\geq 0\}$ (the medial axis of the manifold $S$) where $\phi$ is not differentiable, and therefore the function $\phi$ is not $\rho$-analytic. It follows from Corollary 6.16 that the solution $\phi$ is $H^{1}$-computable in at least super-polynomial time, establishing the claim. ∎
@@ 8 Conclusion
In this paper, we introduced a variational framework for analyzing the computational complexity of partial differential equations (PDEs) without requiring closed-form solutions, thereby generalizing results from previous works [bacho2023complexityblowupsolutionslaplace, Pour-El1983-ls]. Specifically, the variational framework allows to estimate the complexity of a discrete gradient flow, approximation the solution operator of the variational problem, and uses the convergence rates of these discretizations to estimate the complexity of the infinite-dimensional solution of the PDE. This approach enabled us to relate structural properties of PDE operators and their associated variational losses to the computational complexity of the corresponding solutions. Moreover, we introduce a method for proving complexity blowup of the solution based on the regularity-preserving properties of the solution operators. Specifically, if the solution loses regularity relative to analytical boundary and/or initial data, then the solution exhibits a complexity blowup. This result gives a direct link between regularity and approximation theory of PDEs and complexity theory.
@@ A Banach space $X$ is said to have a computability structure, if the following axioms are fulfilled:
One way to prove the polynomial computability of the Legendre grid and weights is to use the Golub–Welsch (GW) algorithm [GW-Method]. For simplicity, consider the case $d=1$. Let $J_{n}\in\mathbb{R}^{(n+1)\times(n+1)}$ be the Jacobi matrix, defined as the tri-diagonal matrix with entries $(J_{n})_{i,j}=0$ for $j\not\in\{i+1,i\}$ for $i\in\{1,\dots,n\}$ and $(J_{n})_{i-1,i}=\beta_{i+1}$$(J_{n})_{i+1,i}=\beta_{i}$ for $i\in\{2,\dots,n\}$, where $\beta_{i}:=\frac{i}{\sqrt{(2i-1)(2i+1)}}$. The Legendre nodes $P_{n}(\Omega)$ are then obtained as the eigenvalues of $J_{n}$. Since the coefficients $\beta_{i}$ are computable in polynomial time, and the tri-diagonal structure allows eigenvalues of $J_{n}$ to be computed in $\mathcal{O}(n^{2})$ operations [Hale2013-zy], this yields the desired complexity bound.
An alternative proof of polynomial computability uses more modern approaches, such as Newton–Raphson iterations with carefully chosen initializations [Petras1999-kp]. These methods achieve improved complexity, requiring only $\mathcal{O}(n)$ [Hale2013-zy] or even $\mathcal{O}(1)$ [Bogaert2012-wf] operations, and currently represent the state of the art for computing Legendre quadratures. ∎
@@
First, observe that the Legendre polynomials can be expressed via the Rodriguez formula [Arfken2013-ny] as derivatives of the polynomial $(x^{2}-1)^{n}$, i.e., the 1-D Legendre polynomial of degree $n\in\mathbb{N}$ is given by
@@
Since $(x^{2}-1)^{n}$ is analytic and $C^{0}$-computable, its derivatives are also $C^{0}$-computable [Pour-El1989-yg] and therefore $H^{k}$-computable. In higher dimensions, the Legendre polynomial $P_{\alpha}\in\Pi_{n}(\Omega)$ for $\alpha\in A_{n,d}$ is given as a finite product of 1-D basis elements. Consequently, the Legendre polynomials themselves are $C^{0}$-computable and therefore $H^{k}$-computable.
Moreover, the roots of these polynomials are computable [Pour-El1989-yg], ensuring that the Legendre grid $P_{n}(\Omega)\subseteq\Omega$ is computable. By the $H^{k}$-computability of $f$, the finite sequence $f(P_{n}(\Omega)):=(f(p_{\alpha}))_{\alpha\in A_{n,d}}\subseteq\mathbb{R}$ is also computable.
@@ Appendix C Appendix
Next, we show the coercivity estimate from (58). Let $u\in H^{1}(\Omega)$ and $\mathcal{E}:H^{1/2}(\partial\Omega)\rightarrow H^{1}(\Omega)$ be the extension operator [Necas2010-yx]. We define $u_{b}:=\mathcal{E}[u|_{\partial\Omega}]$ and $u_{0}:=u-u_{b}$. Then, by Cauchy-Schwarz, the triangle inequality and the extension theorem [HAJLASZ1997221], the following holds
@@
Therefore, applying the estimates in (99) and 95, we have that
@@
To finish the proof we apply Tonelli’s direct method [Maso1993] to a convex, coercive and lower-semi-continuous functional, yielding the existence of the solution. The uniqueness holds since $\mathcal{L}$ is $\lambda$-convex and therefore strictly convex. ∎